OpenStudy (batman4ever):

one more please. the length of a rectangle is 3 feet longer than its width. determine the length and width of the rectangle if its area is 40 square feet.

OpenStudy (watchmath):

L=W+3 A=L*W=(W+3)W=W^2+3W So W^2+3W=40 W^2+3W-40=0 (W+8)(W-5)=0 W=-8 or W=5 Since W is positive, then W=5 Then L=W+3=5+3=8.

OpenStudy (batman4ever):

yay!!!! thank you :)

OpenStudy (watchmath):

Hit the good answer to give me a medal if you like :D